OpenStudy (anonymous):

Suppose an equilateral triangle has sides given by 3y - 4 5y - 18 2y + 3 What is the length of each side of the triangle?

OpenStudy (anonymous):

Length of each side is 17. 3y-4=5y-18 {equilateral triangle's sides} 2y=14 y=7 3(7)-4=17 5(7)-18=17 2(7)+3=17 17 is the length of each side of the equilateral triangle.